Suppose that the line y^=5+2x is fitted to the data points (-1,2), (0,7), and (5,15). Determine the sum of the squared residuals.

Sum of the Squared Residuals =
Answer #1

Points y^ = 5 + 2X . ( y - y^)2

( -1 ,2). 3 (2-3)2 = 1

(0, 7) 5 (7- 5)2 = 4

(5,15) 15 (15 - 15)2 = 0

  

Sum of squared residuals = \sum ( y​​​​​​i - yi^)2

Sum of squared residuals = 5